OpenStudy (anonymous):

Find the complex zeros of f: f(x)=-x^2+x-1 Find the complex zeros of f: f(x)=-x^2+x-1 @Mathematics

OpenStudy (agreene):

\[x= \sqrt[3]{-1};x=-(-1)^{2/3}\]

OpenStudy (anonymous):

How is that?

OpenStudy (anonymous):

quadratic formula

OpenStudy (anonymous):

How do you get that answer?

OpenStudy (anonymous):

a = -1, b = 1, c = -1 x = (-1 +- sqrt(-3))/2 = (-1 +- i*sqrt3)/2

OpenStudy (anonymous):

whoops, it's supposed to be -2 on the bottom :)

OpenStudy (lalaly):

\[x=\frac{ -1 \pm \sqrt {1-4(-1)(-1)}}{2(-1)} = \frac{-1 \pm \sqrt{-3}}{-2}\]\[=\frac{1}{2} \pm \frac{\sqrt{3} i}{-2}\]